In an aggregate expenditure model with no government or foreign sectors, represented by C = a + bY and I (an autonomous amount), an increase in the marginal propensity to save causes the multiplier to rise.
Explain in detail: TRUE, FALSE, or UNCERTAIN.
1
Expert's answer
2015-02-24T09:27:50-0500
In an aggregate expenditure model with no government or foreign sectors, an increase in the marginal propensity to save causes the multiplier to decrease, because the formula for calculating multiplier is m = 1/(1 - c) = 1/s, where s - marginal propensity to save.
So, the statement is False.